2/3/4/16/12 3/4/5/29/20 4/5/6/XX/30 5/6/7/67/42 6/7/8/92/56 find x?

2/3/4/16/12

3/4/5/29/20

4/5/6/XX/30

5/6/7/67/42

6/7/8/92/56
find x?

Grade:upto college level

4 Answers

Saurabh Singh
10 years ago
Interesting series!

first three numbers in each term are pretty simple. they are consecutive numbers

the last number is the product of second and third number
ex.
4*3=12
4*5=20
6*7=42
7*8=56

the second last number is the last number plus the square of first number
ex.
12+2^2=16
20+3^2=29
42+5^2=67
56+6^2=92

hence following the same pattern
30+4^2=XX=46

Thanks & Regards

Saurabh Singh,

askIITians Faculty

B.Tech.

IIT Kanpur

san
10 years ago
sir i did this ques by another way but it took me about 5 10 mins to think this is wht i thought: 16+13=29; 29+17=46; 46+21=67; 67+25=92; i never looked at your answer.. i jus wanted to know that an iitian would have solved this ques at one go? .. if yes then should work even more hard to clear iit
san
10 years ago
sir i did this ques by another way but it took me about 5 10 mins to think this is wht i thought: 16+13=29; 29+17=46; 46+21=67; 67+25=92; i never looked at your answer.. i jus wanted to know that an iitian would have solved this ques at one go? .. if yes then should i work even more hard to clear iit
naina
9 years ago
i think answer shuld be like this allso
4×3+4=16
5×4+5+4=29
6×5+6+5+4=45
7×6+7+6+5+7=67
8×7+8+7+6+8+7=92